# The Challenge of Using LLMs to Simulate Human Behavior: A Causal Inference Perspective

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive potential to simulate human behavior. We identify a fundamental challenge in using them to simulate experiments: when LLM-simulated subjects are blind to the experimental design (as is standard practice with human subjects), variations in treatment systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption. Using demand estimation as a context and an actual experiment as a benchmark, we show this can lead to implausible results. While confounding may in principle be addressed by controlling for covariates, this can compromise ecological validity in the context of LLM simulations: controlled covariates become artificially salient in the simulated decision process, which introduces focalism. This trade-off between unconfoundedness and ecological validity is usually absent in traditional experimental design and represents a unique challenge in LLM simulations. We formalize this challenge theoretically, showing it stems from ambiguous prompting strategies, and hence cannot be fully addressed by improving training data or by fine-tuning. Alternative approaches that unblind the experimental design to the LLM show promise. Our findings suggest that effectively leveraging LLMs for experimental simulations requires fundamentally rethinking established experimental design practices rather than simply adapting protocols developed for human subjects.
@@ 1 Introduction
The rise of large language models (LLMs) like GPT has sparked interest across disciplines (including marketing, computer science, economics, psychology and political science) in leveraging these tools to simulate how humans would respond to questions or stimuli in different contexts (Aher et al.,, 2023; Argyle et al.,, 2023; Arora et al.,, 2024; Brand et al.,, 2023; Dillion et al.,, 2023; Goli and Singh,, 2024; Horton,, 2023; Park et al.,, 2023). If these LLM-simulated experiments accurately approximate real human behavior, the implications for both academics and practitioners are substantial. Academics could use LLMs for pilot experiments to pinpoint stimuli with significant impact, thus improving the efficiency of theory development and experimental design. In fact, several review articles have already been written about this recent yet fast-growing literature (Bail,, 2024; Hutson and Mastin,, 2023; Sarstedt et al.,, 2024; Simmons and Hare,, 2023). Firms could leverage these realistic simulations to explore different ideas and strategies, thereby improving customer insight and product development and optimizing marketing mix variables. Accordingly, in the recent past we have witnessed a large influx of firms offering services leveraging LLMs for customer insights (e.g., Synthetic Users, Outset AI, Nexxt, Voxpopme, Evidenza, Expected Parrot, Meaningful, xPolls, Ipsos, and CivicSync).
@@
Many practical applications of LLM simulations involve exploring counterfactual scenarios where one factor is altered while others remain constant, hence requiring causal estimates of treatment effects. For example, what would happen to sales if prices were increased, how would consumers respond if a message were framed differently, or how would market dynamics shift with the introduction of a new product? While the success of using LLMs to simulate counterfactual depends on the size and quality of their training or validation data that can correct for potential bias (Ludwig et al.,, 2024; Wang et al.,, 2024), the success also critically depends on the design of the simulations.
Although extensive literature on experimental design with human subjects has established best practices and universal standards—such as random assignment and blinding—it remains unclear what constitutes best practices in designing LLM simulations and how these might differ from traditional experimental design principles. Establishing LLM experimental design principles is crucial for enhancing the reliability and validity of LLM simulation results, thereby making them more valuable for both researchers and practitioners. Without clear guidelines on which experimental designs to adopt or avoid, the vast flexibility and sensitivity in LLM simulations can lead to widely varying outcomes (Carlson and Burbano,, 2024; Gao et al.,, 2024; Ludwig et al.,, 2024; Mohammadi,, 2024). When different simulation approaches yield conflicting results, researchers lack a basis for determining how much weight to place on each finding. By developing standardized design principles, researchers can assign greater weight to well-designed simulations, enhancing the reliability and usefulness of the results. Developing such design principle is essential for advancing the field and ensuring that LLM simulations effectively contribute to our understanding of human behavior.
@@
Many LLM simulations to date have relied on blinded experimental designs, where simulated participants are not aware of the experimental conditions or hypotheses being tested (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023). Blinding [^1] Also sometimes referred to as “masking” (Association et al.,, 2019). has many benefits in human subject research: it reduces demand effects where participants modify their behavior based on what they believe researchers want to see and often helps maintain ecological validity by presenting scenarios that more closely mirror real-world decision contexts. Following this convention, researchers typically construct LLM prompts that present a single scenario without revealing its experimental nature, similar to how human participants would experience the treatment condition in a traditional between-subjects design.
@@
Although blinding is considered a best practice in traditional experiments, we show that it can have unintended consequences in LLM simulations which may no longer justify its benefits. Unlike human subjects and contexts that exist independently of the experiment, LLM-simulated subjects and contexts are ”created on the fly” based on the prompt. When researchers make LLM-simulated “subjects” blind to experimental conditions, variations in the treatment can systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption typically required for valid causal inference.
@@
We use demand estimation as an example to empirically illustrate this challenge. We conduct experiments simulating consumer demand at different price points. We find that altering the product’s price not only affects the purchase outcomes but also systematically influences other variables that should not be influenced by the treatment—such as pre-treatment variables (e.g., demographic characteristics and customer purchase history) and contextual factors (e.g., price of competing products). This unintended variation in unspecified covariates violates the unconfoundedness assumptions, and we show that it also generates implausible demand curves when compared against an actual experiment with a representative sample of human participants.
@@
A tempting solution is specifying more variables in the prompts to address this confounding issue. However, we find this approach may sacrifice ecological validity because including more covariates can make the scenarios presented to the LLM less reflective of real-world decision-making contexts. For instance, when competing product information is explicitly mentioned in prompts, these variables become artificially salient in the LLM’s decision process, unlike typical consumers who may not always consider such information during routine purchases. This focalism issue creates a fundamental tension between controlling for confounding factors and maintaining ecological validity - a tradeoff that typically does not exist in traditional experiments where covariates can be measured and controlled for during data analysis without making them artificially salient to participants during data generation.
@@
We present a theoretical framework illustrating that confoundedness in blind LLM simulations stems from prompts being ambiguous. To overcome this fundamental limitation and improve the simulation results, researchers should consider relaxing the blinding requirement and informing the LLM of the experimental design. This approach is not subject to the same confounding issues and holds promise for better generalizable performance as LLMs continue to advance. To complement this approach, an opportunity for future research is to develop LLMs that can better interpret and understand experimental designs as well as natural language descriptions of counterfactual scenarios.
More broadly, we aim to contribute to the growing literature on LLM simulations by providing a starting point for establishing best practices in their design. By highlighting the trade-offs between blinding, unconfoundedness, and ecological validity, we offer insights that can help researchers and practitioners design more reliable and valid LLM-based experiments. Our work underscores the importance of adjusting traditional experimental design principles to the unique characteristics of LLMs, ensuring that these powerful tools can be effectively harnessed to simulate and understand human behavior.
@@
The rest of the paper is organized as follows. Section 2 demonstrates the unintended consequences of blinding empirically using demand estimation as an example. Section 3 explores the standard solution of controlling for covariates and highlights its unique challenges in the context of LLM simulations. Section 4 proves an impossibility theorem demonstrating this fundamental issue will persist for future models as long as ambiguous prompting strategies are used, and examines its implications for techniques like fine-tuning and RAG. Section 5 discusses promising directions for making prompts unambiguous through unblinding, and Section 6 concludes.
@@ 2 Unintended Confounding in Blind LLM Simulations
One common way to implement blinding in LLM simulations is to conduct between-prompt experiments, where each prompt represents a single experimental condition without revealing the broader experimental context (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023). Blinded between-prompt LLM experiments have been used to mimic both between-subject and within-subject experiments with human participants. For example, to mimic a between-subject experiment that studies price sensitivity, researchers might present different prices to the LLM in separate prompts, with each prompt describing only the current price scenario without mentioning the experimental nature of the price variation. To mimic a within-subject conjoint analysis study, researchers might present different product profiles or choice sets to the LLM in separate prompts, with each prompt describing only the current product profile or choice set.
@@
While random assignment in traditional human experiments ensures pre-treatment variables are balanced across conditions and contextual variables remain constant (even when unobserved), LLM-simulated participants and environments are generated dynamically. When researchers leave details unspecified in blind LLM prompts, the model must make (implicit or explicit) assumptions about these unspecified variables, which can systematically vary with the treatment and introduce confounding factors.
@@
We illustrate these issues using demand estimation, a classical problem in economics and marketing, where researchers are often interested in understanding how changes in a product’s price affect demand while holding everything else constant. We select the top 40 Consumer Packaged Goods (CPG) categories according to DellaVigna and Gentzkow, (2019), and for each category we pick one of the top selling products according to Walmart.com. We record the regular price of each product on Walmart.com around April, 2024. Table A.1 in the Appendix documents this set of products. For each product, we vary the price from 0 to 200% of the regular price, in 20% increments. This gives us 11 price points to test for each product ($0\%\}$ , 20 % , ..100 % , ..200 % } of the regular price). Throughout the paper, for each prompt template and for each {category,product} pair described in Table A.1, we vary the price systematically across these 11 price points. We refer to “relative price” as the price relative to the regular price ($\{-1,-0.8,...,0,...,0.8,1\}$ where 00 is the regular price). We use the Open AI API and we “pin” the model to gpt-4o-mini-2024-07-18 (the latest available at the time of running the experiment) to make sure that the exact same model is used for all our simulations. We adjust the “System” part of the prompt to increase the probability of receiving a usable answer from the model (the “System” portion of the prompt, which can be manipulated on the API and which is set to “You are a helpful assistant” on ChatGPT, provides meta-instructions to GPT). Unless noted otherwise, for each prompt that corresponds to a {category, product, price} combination, we set the temperature to 1 and generate 50 draws to calculate the average response.
@@
First, we elicit unspecified variables from the LLM. For each {category, product, price} combination, we submit a prompt that asks GPT to fill in the blank contextual or pre-treatment variables given the randomized price (Kurita et al.,, 2019; Nozza et al.,, 2022; Zhang et al.,, 2020). Prompt 2 gives an example of such template that asks the past price given the current randomized price.
@@
Prompt 1: Unintended correlation between price and past price System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The last time you purchased this product, it was priced at \$          [a number with up to 2 decimal points]. The product is currently priced at: {price}. Return example 1: XX.XX
@@
Figure 1 demonstrates positive correlations between the price of the focal product and several variables, including the price that the customer paid last time, the price of the competing product, and the expiration days of the product. The x-axis is price relative to the baseline price, allowing us to aggregate across products.333 Prompts B and B in Appendix B document the prompts used to elicit the correlations between randomized price and competing price/expiration date. Appendix C presents additional analysis demonstrating the broader scope of these unintended correlations across multiple variables. While these correlations are plausible in observational data, they are undesirable in an experiment where researchers are interested in understanding what happens when estimating the treatment effect of changing the price of the focal product while holding everything else constant.
@@
Next, we explore whether these unintended correlations are associated with confounding and biased estimates of relevant quantities when the goal is estimating causal treatment effects. To that end, we run an actual survey on human subjects, use a similar prompt in an LLM simulation, and compare the results. The survey with human participants (pre-registered, see https://aspredicted.org/pqys-st6w.pdf) was developed using Qualtrics. Each respondent answered one purchase intention question per product, with prices randomly drawn for each product and the order of products randomized across respondent. [^4] We acknowledge that one difference between the study run with humans vs. GPT is that each human answered 40 questions, while on GPT a different independent call to the API was made for each question (following standard practice). Following our pre-registration, we collected data from 1,000 respondents, and excluded 9 respondents according to pre-registered exclusion criteria (based on attention filters at the beginning of the survey). The respondents were recruited from Prolific, where we paid a premium to obtain a representative sample of the US population (based on census data). We then use Prompt 2 to simulate potential purchase decisions at different price points (where again {category,product} and {price} were varied according to our experimental design, mirroring the experiment with humans), and for each corresponding prompt, we used gpt-4o-mini-2024-07-18, with temperature set to 1, to generate 50 draws to calculate the average purchase probability at each {category,product,price} combination. Figure 2 shows the average demand curves obtained from GPT and from human participants. While the curves elicited from human participants follow a classic downward-sloping demand curve, the GPT-simulated curves are different from the curves elicited from human participants and follow an inverted-U shape. Although it is understandable that it is challenging for GPT to predict demand when price equals 0 as such scenario is probably not well represented in its training data, the curve remains inverted-U shape even if we remove the first point corresponding to a price of zero.
@@
Figure 2: Demand curve elicited from humans vs. LLM
These findings are not surprising in hindsight: in real observational data, the price of the focal product should be informative of the price of the competing products, especially if we do not know anything about the context, the type of store, and how the price for a product is set. An LLM trained on such data would also naturally make such association. This association can be realistic and useful if researchers intend to use LLMs to simulate competing prices based on the focal product price. However, this same association is problematic when researchers intend to estimate the causal impact of changing the focal price, while holding everything else constant, because it leads to endogeneity. One common approach to addressing such endogeneity is to control for additional covariates. The next section explores the benefits and novel challenges associated with this approach.

### v3

@@ Abstract
Large Language Models (LLMs) have shown impressive potential to simulate human behavior. We identify a fundamental challenge in using them to simulate experiments: when LLM-simulated subjects are blind to the experimental design (as is standard practice with human subjects), variations in treatment systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption. Using demand estimation as a context and an actual experiment with 40 different products as a benchmark, we show this can lead to implausible results. While confounding may in principle be addressed by controlling for covariates, this can compromise ecological validity in the context of LLM simulations: controlled covariates become artificially salient in the simulated decision process. We show formally that confoundness stems from ambiguous prompting strategies. Therefore, it can be addressed by developing unambiguous prompting strategies through unblinding, i.e., revealing the experiment design in LLM simulations. Our empirical results show that this strategy consistently enhances model performance across all tested models, including both out-of-box reasoning and non-reasoning models. We also show that it is a technique that complements fine-tuning: while fine-tuning can improve simulation performance, an unambiguous prompting strategy makes the predictions robust to the inclusion of irrelevant data in the fine-tuning process.
@@ 1 Introduction
The rise of large language models (LLMs) like GPT has sparked interest across disciplines (including marketing, computer science, economics, psychology and political science) in leveraging these tools to simulate how humans would respond to questions or stimuli in different contexts (aher_using_2023; argyle_out_2023; arora2024express; binz2025foundation; dillion_can_2023; goli2024frontiers; hewitt2024predicting; horton_large_2023; park_generative_2023; qin2024aiturk; toubia2025twin). If these LLM-simulated experiments accurately approximate real human behavior, the implications for both academics and practitioners are substantial. Academics could use LLMs for pilot experiments to pinpoint stimuli with significant impact, thus improving the efficiency of theory development and experimental design. In fact, several review articles have already been written about this recent yet fast-growing literature (bail2024can; hutson2023guinea; sarstedt2024using; simmons2023large). Firms could leverage these realistic simulations to explore different ideas and strategies, thereby improving customer insight and product development and optimizing marketing mix variables. Accordingly, in the recent past we have witnessed a large influx of firms offering services leveraging LLMs for customer insights (e.g., Synthetic Users, Outset AI, Nexxt, Voxpopme, Evidenza, Expected Parrot, Meaningful, xPolls, Ipsos, Pyxis and Generation Lab).
@@
In this paper we focus specifically on leveraging LLMs to simulate counterfactual human behavior. Indeed, many practical applications of LLM simulations involve exploring counterfactual scenarios where one factor is altered while others remain constant. For example, what would happen to sales if prices were increased, how would consumers respond if a message were framed differently, or how would market dynamics shift with the introduction of a new product?
Early evidence on LLMs’ ability to simulate counterfactual human behavior is encouraging but mixed. For example, across 70 nationally-representative U.S. studies, hewitt2024predicting find that the correlation between predicted and observed treatment effect size is large ($r=0.85$), but the raw predictions derived from GPT-4 systematically overestimated actual effects, leading to a large (relative to human forecasters) RMSE between predicted and observed effects. Across 11 behavioral economics experiments, toubia2025twin find that LLM simulations achieve high predictive accuracy at the individual level, but replicated only about half of the treatment effects observed at the aggregate level.
The ability of LLMs to simulate counterfactuals will likely improve as more powerful models are developed and the size and quality of the training and validation data continues to increase (ludwig2024large; wang2024large). In this paper, we focus on another key factor, which may be in fact much easier and cheaper to address and which is under the control of any user: the design of LLM experiments. We identify a fundamental issue that arises when simply applying best practices from human experiments to LLM experiments. This issue is not specific to a particular model or training dataset and, if left unaddressed, it is likely to cause some amount of residual bias in LLM simulations even as accuracy improves with better models and data. At the same time, we find this issue can be alleviated by carefully adjusting experimental best practices to the world of LLMs.
@@
In particular, many LLM simulations to date have relied on blinded experimental designs, where simulated participants are not aware of the experimental conditions or hypotheses being tested (argyle_out_2023; brand_using_2023; hewitt2024predicting; horton_large_2023; toubia2025twin). Blinding [^1] Also sometimes referred to as “masking” (american2019publication). has many benefits in human subject research: it reduces demand effects where participants modify their behavior based on what they believe researchers want to see, and it helps maintain ecological validity by presenting scenarios that more closely mirror real-world decision contexts. Following this convention, researchers typically construct LLM prompts that present a single scenario without revealing its experimental nature, similar to how human participants would experience the treatment condition in a between-subjects design.
@@
Although blinding is considered a best practice in human-subject experiments, we show that it can have unintended consequences in LLM simulations, which may no longer justify its benefits. Unlike human subjects and contexts that exist independently of the experiment, LLM-simulated subjects and contexts are ”created on the fly” based on prompts. When researchers make LLM-simulated “subjects” blind to experimental conditions, variations in the treatment can systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption typically required for valid causal inference.
@@
We use demand estimation as an example to empirically illustrate this challenge. We conduct experiments simulating consumer demand for 40 different products at different price points. We find that altering the product’s price not only affects the purchase outcomes but also systematically influences other variables that should not be influenced by the treatment—such as pre-treatment variables (e.g., demographic characteristics and customer purchase history) and contextual factors (e.g., price of competing products). This unintended variation in unspecified covariates violates the unconfoundedness assumptions, and in our case it also generates implausible demand curves when compared against an actual experiment with a representative sample of human participants.
@@
A tempting solution is specifying more variables in the prompts to address this confounding issue. However, we find this approach may sacrifice ecological validity because including more covariates can make the scenarios presented to the LLM less reflective of real-world decision-making contexts. For instance, when competing product information is explicitly mentioned in prompts, these variables become artificially salient in the LLM’s decision process, unlike typical consumers who may not always consider such information during routine purchases. This focalism issue creates a fundamental tension between controlling for confounding factors and maintaining ecological validity - a tradeoff that typically does not exist in human-subject experiments where covariates can be measured and controlled for during data analysis without making them artificially salient to participants during data generation.
@@
We present a theoretical framework illustrating that confoundedness in blind LLM simulations stems from prompts being ambiguous. To overcome this fundamental limitation and improve the simulation results, researchers should consider relaxing the blinding requirement and informing the LLM of the experimental design. This approach is not subject to the same confounding issues and holds promise for better generalizable performance as LLMs continue to advance.
More broadly, we aim to contribute to the growing literature on LLM simulations by helping establish best practices in their design. Although extensive literature on experimental design with human subjects has established best practices and universal standards—such as random assignment and blinding—it remains unclear what constitutes best practices in designing LLM simulations and how these might differ from human-subject experimental design principles. Establishing LLM experimental design principles is crucial for enhancing the reliability and validity of LLM simulation results, thereby making them more valuable for both researchers and practitioners. Without clear guidelines on which experimental designs to adopt or avoid, the vast flexibility and sensitivity in LLM simulations can lead to widely varying outcomes (carlson2024use; gao2024take; ludwig2024large; mohammadi2024wait). When different simulation approaches yield conflicting results, researchers lack a basis for determining how much weight to place on each finding. By developing standardized design principles, researchers can assign greater weight to well-designed simulations, enhancing the reliability and usefulness of the results. Developing such design principle is essential for advancing the field and ensuring that LLM simulations effectively contribute to our understanding of human behavior. In particular, by highlighting the trade-offs between blinding, unconfoundedness, and ecological validity, we offer insights that can help researchers and practitioners design more reliable and valid LLM-based experiments. In doing so, our work underscores the importance of adjusting human-subject experimental design principles to the unique characteristics of LLMs, ensuring that these powerful tools can be effectively harnessed to simulate and understand human behavior.
@@
The rest of the paper is organized as follows. Section 2 empirically demonstrates the unintended consequences of blinding using demand estimation as an example. Section 3 explores the standard solution of controlling for covariates and highlights its unique challenges in the context of LLM simulations. Section 4 argues theoretically that this fundamental issue will persist for future models as long as ambiguous prompting strategies are used. Section 5 discusses promising directions for making prompts unambiguous through unblinding, and Section 6 concludes.
@@ 2 Unintended Confounding in Blind LLM Simulations
One common way to implement blinding in LLM simulations is to conduct between-prompt experiments, where each prompt represents a single experimental condition without revealing the broader experimental context (argyle_out_2023; brand_using_2023; hewitt2024predicting; horton_large_2023; toubia2025twin). Blinded between-prompt LLM experiments have been used to mimic both between-subject and within-subject experiments with human participants. For example, to mimic a between-subject experiment that studies price sensitivity, researchers might present different prices to the LLM in separate prompts, with each prompt describing only the current price scenario without mentioning the experimental nature of the price variation. To mimic a within-subject conjoint analysis study, researchers might present different product profiles or choice sets to the LLM in separate prompts, with each prompt describing only the current product profile or choice set.
@@
While random assignment in human-subject experiments ensures pre-treatment variables are balanced across conditions and contextual variables remain constant (even when unobserved), LLM-simulated participants and environments are generated dynamically based on prompts. When researchers leave details unspecified in blind LLM prompts, the model may make (implicit or explicit) assumptions about these unspecified variables, which can systematically vary with the treatment and introduce confounding factors.
@@
We illustrate these issues using demand estimation, a classical problem in economics and marketing, where researchers are often interested in understanding how changes in a product’s price affect demand while holding everything else constant. We select the top 40 Consumer Packaged Goods (CPG) categories according to dellavigna_uniform_2019, and for each category we pick one of the top selling products according to Walmart.com. We record the regular price of each product on Walmart.com around April, 2024. Table A.1 in the Web Appendix documents this set of 40 products. For each product, we vary the price from 0 to 200% of the regular price, in 20% increments. This gives us 11 price points to test for each product ($\{0,20\%,..100\%,..200\%\}$ of the regular price). Throughout the paper, for each prompt template and for each {category,product} pair described in Table A.1, we vary the price systematically across these 11 price points. We refer to “relative price” as the price relative to the regular price ($\{-1,-0.8,...,0,...,0.8,1\}$ where 0 is the regular price). We use the Open AI API and we “pin” the model to gpt-4o-mini-2024-07-18 (the latest available at the time of running the experiment) to make sure that the exact same model is used for all our simulations. We adjust the “System” part of the prompt to increase the probability of receiving a usable answer from the model (the “System” portion of the prompt, which can be manipulated on the API, provides meta-instructions to GPT). Unless noted otherwise, for each prompt that corresponds to a {category, product, price} combination, we set the temperature to 1 and generate 50 draws to calculate the average response.
@@
First, we elicit unspecified variables from the LLM. For each {category, product, price} combination, we submit a prompt that asks GPT to fill in the blank contextual or pre-treatment variables given the randomized price (kurita2019measuring; nozza2022pipelines; zhang2020hurtful). Prompt 2 gives an example of such template that asks the past price given the current randomized price.
@@
Prompt 1: Unintended correlation between price and past price System: You, AI, are a customer. Your task is to fill in the blanks $_$. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The last time you purchased this product, it was priced at \$     [a number with up to 2 decimal points]. The product is currently priced at: {price}. Return example 1: $X$.XX
@@
Figure 1 demonstrates positive correlations between the price of the focal product and several variables, including the price that the customer paid last time, the price of the competing product, and the expiration days of the product. The x-axis is price relative to the baseline price, allowing us to aggregate across products. [^3] Prompts B and B in Web Appendix B document the prompts used to elicit the correlations between randomized price and competing price/expiration date. Web Appendix C presents additional analysis demonstrating the broader scope of these unintended correlations across multiple variables. While these correlations are plausible in observational data, they are undesirable in an experiment where researchers are interested in understanding what happens when changing the price of the focal product while holding everything else constant.
@@
Next, we explore whether these unintended correlations are associated with biased estimates of relevant quantities when the goal is simulate counterfactual human behavior. We run an actual survey on human subjects, using a similar prompt in an LLM simulation, and compare the results. The survey with human participants (pre-registered, see https://aspredicted.org/pqys-st6w.pdf) was developed using Qualtrics. Each respondent answered one purchase intention question for each of the 40 products (see Appendix A for the complete list of categories, products, and prices), with prices randomly drawn for each product and the order of products randomized across respondent. [^4] We acknowledge that one difference between the study run with humans vs. GPT is that each human answered 40 questions, while on GPT a different independent call to the API was made for each question (following standard practice). Following our pre-registration, we collected data from 1,000 respondents, and excluded 9 respondents according to pre-registered exclusion criteria (based on attention filters at the beginning of the survey). The respondents were recruited from Prolific, where we paid a premium to obtain a representative sample of the US population (based on census data). We then use Prompt 2 to simulate potential purchase decisions at different price points (where again {category,product} and {price} were varied according to our experimental design, mirroring the experiment with humans), and for each corresponding prompt, we used gpt-4o-mini-2024-07-18, with temperature set to 1, to generate 50 draws to calculate the average purchase probability at each {category,product,price} combination. Figure 2 shows the average demand curves (averaged over the 40 products) obtained from GPT and from human participants. While the curves elicited from human participants follow a classic downward-sloping demand pattern, the GPT-simulated curves follow an inverted-U shape. Although it is understandable that it is challenging for GPT to predict demand when price equals 0 as such scenario is probably not well represented in its training data, the curve remains inverted-U shaped even if we remove the first point corresponding to a price of zero.
@@
Figure 2: Demand curve elicited from humans vs. LLM, averaged over 40 products
These findings are not surprising in hindsight: in real observational data, the price of the focal product should be informative of other factors such as the price of the competing products, especially if we do not know anything about the context, the type of store, and how the price for a product is set. An LLM trained on such data would also naturally make such association. This association can be realistic and useful if researchers intend to use LLMs to elicit the correlation between these other factors and the focal product price. However, this same association is problematic when researchers intend to estimate the causal impact of changing the focal price, while holding everything else constant, because it leads to endogeneity. One common approach to addressing such endogeneity is to control for additional covariates. The next section explores the benefits and novel challenges associated with this approach.